What is the final velocity of a 1200 kilogram car starting from rest after 50 meters when 4500 Newton's of force is applied?


1
Expert's answer
2021-05-04T12:09:10-0400

According to the second Newton's law, the acceleration of the car is:


"a = \dfrac{F}{m}"

where "F = 4500N" is the total force applied to the car and "m = 1200kg" is the mass of the car.

According to the kinematic equation, the distance covered under a constant acceleration is:


"d =v_0t + \dfrac{at^2}{2} = \dfrac{at^2}{2}"

where "v_0 = 0" is the initial speed of the car (equal to 0 since the car starts from rest) and "t" is the time.

Since "d = 50m", we can find the time:


"t = \sqrt{\dfrac{2d}{a}} = \sqrt{\dfrac{2dm}{F}} \\"

The velocity of the car at this moment of time (if it starts from rest) is given by the following kinematic equation:


"v =v_0 + at = at"

Substituting the expressions for "a" and "t", obtain:


"v = \dfrac{F}{m}\cdot \sqrt{\dfrac{2dm}{F}} = \sqrt{\dfrac{2dF}{m}}\\
v = \sqrt{\dfrac{2\cdot 50\cdot 4500}{1200}} \approx 6.1 m/s"

Answer. 6.1 m/s.